OpenStudy (hexagon001):

I dont get this... all comes ahead... its all in the context of evaluating improper integrals in complex analysis..

OpenStudy (hexagon001):

\[\int\limits_{-\infty}^{\infty}sint dt=\lim_{r \rightarrow \infty}\int\limits_{-r}^{r}sint dt =\lim_{r \rightarrow \infty}[-cost]=\lim_{r \rightarrow \infty}0=0\] note; the value of [−cost] is between r and -r. so my question is how have they got that the limit of -cost is 0?

OpenStudy (hexagon001):

we dont even know the values of r and -r...so how did htey find that -cost has limit 0?

terenzreignz (terenzreignz):

I'm taking this at face-value, for now, but what if we forget about the limit for the moment, and focus on the integral \[\huge \int\limits_{-r}^{r}\sin \ t \ dt\]instead?

OpenStudy (experimentx):

use the fact that cos(-x) = cos(x) ... this integral shouldn't converge (i think). what you are calculating is Cauchy Principal Value.

terenzreignz (terenzreignz):

\[\large = \left[-\cos \ t \right]_{-r}^{r}=(-\cos \ r)-(-\cos(-r))=0\]

OpenStudy (agent0smith):

Bit too big, terenz :P But yes this integral should not converge... Area under the sine curve will never converge... but this may have something to do with the "...in complex analysis" :/

terenzreignz (terenzreignz):

Oh, this was complex analysis? That's beyond my knowledge. Take it away, guys :D

OpenStudy (agent0smith):

lol i have no idea, I don't remember studying integrals of complex numbers/functions in university calculus.

OpenStudy (hexagon001):

ahaaa moment of epiphany..thanks terenzreignz

terenzreignz (terenzreignz):

All in a day's work? :D

OpenStudy (hexagon001):

lol

OpenStudy (agent0smith):

But that still doesn't change the fact that the area under a sine/cosine curve is unbounded...

terenzreignz (terenzreignz):

Yeah... that's why I was careful and said I'd take it at face value... whatever that means, LOL

OpenStudy (experimentx):

yea this doesn't converge. this is called http://en.wikipedia.org/wiki/Cauchy_principal_value

OpenStudy (anonymous):

I thought that's because the total area of under the sine curve on the right is equal to the area on the left and hence the can cancel out\[\int\limits^r_{-r}\]

OpenStudy (agent0smith):

^yeah, I'm guessing that's what it's getting at. Idk anything about the cauchy principal value thought.

OpenStudy (experimentx):

no they don't ... both are indeterminate.

OpenStudy (hexagon001):

Honestly...wish complex numbers did not exist...they do really make things more.....complex.. :D

OpenStudy (anonymous):

I do think complex number makes life easier. :P

OpenStudy (experimentx):

^^ agree with agostino

OpenStudy (hexagon001):

i guess I do agree to an extent but all the studying.... blurgh...

OpenStudy (experimentx):

yeah yeah ... currently I am also stuck on complex analysis.

OpenStudy (hexagon001):

I feel for you..

OpenStudy (experimentx):

what book are you following?

OpenStudy (hexagon001):

I'm studying from a university published book-the open univeristy

OpenStudy (experimentx):

i don't know what it is.

OpenStudy (hexagon001):

I'm studying thorugh distant learning/home-schooling.. In the UK we have a univeristy called the Open University which caters for distant learning students...so i am doing my maths degree with them..and they have provided all the materials of study

OpenStudy (experimentx):

Oh .. i know that OU.

OpenStudy (hexagon001):

:) So what book do you use?

OpenStudy (experimentx):

http://books.google.com/books/about/Complex_Variables.html?id=9zKl4lXEXlsC&redir_esc=y currently looking this

OpenStudy (anonymous):

Anyone uses Springer's books?

OpenStudy (hexagon001):

@experimentX just had a look through an online version of the book... imho it is better explained and set out than the book im studying from... thanks for introducing..will definitely be using it. @agostino never used it.